OpenStudy (anonymous):

At the starting gun, a runner accelerates at 1.8m/s2 for 4.7s . The runner's acceleration is zero for the rest of the race. What is the speed of the runner at t=2.0s? What is the speed of the runner at the end of the race?

OpenStudy (anonymous):

\[v_f=v_i+a \times t\] \[v_f = 0+1.8 \times 2 = 3.6_{m/s}\] \[v_{f2} = 0+1.8 \times 4.7 = 8.46_{m/s}\]

OpenStudy (anonymous):

His speed at the end of the race will be the same as the one he gained from accelerating. Acceleration that equals zero = constant speed.